OpenStudy (anonymous):

Hi all! I have a question left on my problem set and I'd like to check my answer :) "verify that the function satisfies the hypotheses of the mean value theorem on the given interval. Then find all numbers c that satisfy the conclusion of the mean value theorem f(x) = 1/x [1,3]"

OpenStudy (anonymous):

So, I have the derivative as f'(x)= -1/(x^2)

OpenStudy (anonymous):

The value c I got is sqrt(3/4)

OpenStudy (anonymous):

I'm a bit confused about this problem :)

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

f(1)=1/1=1 f(3)=1/3 The average slope between these two points \[\frac{ \frac{ 1 }{ 3 } -1 }{ 3-1 }= -\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

f'(x)= -1/(x^2) as you stated before make it equal to -1/3 -1/(x^2)=-1/3 \[x=\pm {\sqrt{3}}\]

OpenStudy (anonymous):

The solution is only the + as the - lies outside the interval [1,3]

OpenStudy (anonymous):

If you need further info check this video: http://www.youtube.com/watch?v=xYOrYLq3fE0